### v4

@@ Appendix C.2  Conditions for Bayes Optimality
In this appendix, we review the result that the maximum likelihood decision rule is Bayes optimal when utility takes the form in (13) and the decision maker’s subjective prior is uniform across realizations of the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}. To show this result, we first establish that the maximum a posteriori decision rule is Bayes optimal for our chosen loss function given any prior. The maximum a posteriori decision rule $f^{*}_{\pi}$ selects the maxima of the posterior distribution of $\boldsymbol{\theta}$}. Formally, let $\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})$ be the set of $\boldsymbol{\theta}$} values that maximize the posterior distribution given the observed data $\boldsymbol{x}$}, i.e.
@@
where $\pi\in\Delta({\boldsymbol{\Theta}})$ is the prior belief about $\boldsymbol{\theta}$}. The decision rule $f^{*}_{\pi}$ can then be defined as follows:
@@
${\mathbb{E}}\big[EU(g,{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(g,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})$
$=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\Bigg[\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})g({\boldsymbol{x}})({\boldsymbol{\theta}})\Bigg]\pi({\boldsymbol{\theta}}).$
By rearranging terms in the summation, we can bound the Bayes expected utility of $g$:
@@
Thus, since $f^{*}_{\pi}$ achieves the upper bound on the Bayes expected utility of any decision rule, we conclude that $f^{*}_{\pi}$ is Bayes optimal. Finally, observe that when the prior distribution $\pi(\theta)$ is constant, the maximizers of the posterior distribution in (15) are simply the maximizers of the likelihood, and the Bayes rule $f^{*}_{\pi}$ in (16) simplifies to our maximum likelihood decision rule.
@@
Appendix D  Construction of Credible Sets
We construct credible sets on the joint distribution of potential outcomes in the sample using the posterior belief implied by a uniform prior and the design-based likelihood. A credible set of level $\alpha$ is a set of joint distributions of potential outcomes covering a posterior probability mass of at least $\alpha$. We report the smallest credible set, which is the credible set containing the fewest points. We construct these sets with the following algorithm:
@@ Appendix D  Construction of Credible Sets
To report credible sets on the number of always takers, compliers, defiers, or never takers individually, we report the range of values for that type present in the credible set constructed on the joint distribution. Note that the credible sets on individual response types are not independent. For example, the maximal values in the individual credible sets for always takers, compliers, defiers, and never takers may not constitute an element of the credible set on the joint distribution of potential outcomes.
Appendix E  Auxiliary Statistics
@@
Figure E.1: Standard Statistics, Proposed Design-Based Maximum Likelihood Estimates, and Auxiliary Statistics

### v5

@@ A Standard Proof Establishes Bayes Optimality:
We first establish that the maximum a posteriori decision rule is Bayes optimal for our chosen loss function given any prior. The maximum a posteriori decision rule $f^{*}_{\pi}$ selects the maxima of the posterior distribution of $\boldsymbol{\theta}$}. Formally, let $\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})$ be the set of $\boldsymbol{\theta}$} values that maximize the posterior distribution given the observed data $\boldsymbol{x}$}, i.e.
@@
where $\pi\in\Delta$ is the prior belief about $\boldsymbol{\theta}$}. The decision rule $f^{*}_{\pi}$ can then be defined as follows:
@@
${\mathbb{E}}\big[EU(g,{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(g,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\Bigg[\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})g({\boldsymbol{x}})({\boldsymbol{\theta}})\Bigg]\pi({\boldsymbol{\theta}}).$
Note that, under our choice of utility function, the expected utility of decision rule $g$ is the probability that $g$ correctly guesses the distribution of potential outcomes $\boldsymbol{\theta}$}. By rearranging terms in the summation, we can bound the Bayes expected utility of $g$:
@@
Thus, since $f^{*}_{\pi}$ achieves the upper bound on the Bayes expected utility of any decision rule, we conclude that $f^{*}_{\pi}$ is Bayes optimal. Finally, observe that when the prior distribution $\pi({\boldsymbol{\theta}})$ is constant, the maximizers of the posterior distribution in (17) are simply the maximizers of the likelihood, and the Bayes rule $f^{*}_{\pi}$ in (18) simplifies to our maximum likelihood decision rule. ∎
@@
The Relative Performance of the Optimal Rule Increases in Even Sample Sizes from 2 to 200:
The Bayes optimality of our proposed maximum likelihood decision rule allows us to quantify the gains from using the optimal rule over alternatives. We consider two alternative decision rules. The first alternative decision rule, which we refer to as the “Fréchet rule,” estimates the marginal distributions of potential outcomes directly according to (4), and then chooses randomly among the joint distributions of potential outcomes in the corresponding Fréchet set. The Fréchet rule $\widehat{{\boldsymbol{\theta}}}_{F}$ satisfies:
The second alternative decision rule, which we refer to as the “monotonicity rule,” approximates the behavior of a statistician who uses the monotonicity assumption of Imbens and Angrist (1994) in the population to estimate the share of compliers or defiers in the sample. The rule uses the observed difference in takeup rate between the intervention and control arms $\delta$ to approximate the share of compliers or defiers. In particular, if the difference is non-negative, the rule imposes a no-defiers monotonicity restriction, and if the difference is non-positive, it imposes a no-compliers monotonicity restriction.
$\hat{\theta}_{10}=n\cdot\max\{\delta,0\},\quad\hat{\theta}_{01}=n\cdot\max\{-\delta,0\},$
$=\bigg(\mathbf{1}_{\{\delta>0\}}\left[n\frac{x_{C1}}{x_{C1}+x_{C0}}\right]+\mathbf{1}_{\{\delta\leq 0\}}\left[n\frac{x_{I1}}{x_{I1}+x_{I0}}\right],\ n\cdot\max\{\delta,0\},\ n\cdot\max\{-\delta,0\},$
$\qquad n-\hat{\theta}_{11}-\hat{\theta}_{01}-\hat{\theta}_{10}\bigg)$
with probability one.
We calculate the exact Bayes expected utility achieved by our optimal maximum likelihood rule and each of the two alternative rules. For a given sample size, we enumerate each of the finitely many feasible sample distributions of potential outcomes $\boldsymbol{\theta}$}. For each distribution $\boldsymbol{\theta}$}, we compute the expected utility of our decision rules by comparing the distribution of estimates they produce under each data realization $\boldsymbol{x}$} to the true value of $\boldsymbol{\theta}$}. We average the resulting utilities over the values of $\boldsymbol{x}$} according to the likelihood. We then average these expected utilities according to a uniform prior for $\boldsymbol{\theta}$} to get the Bayes expected utility.
Figure B.2 plots the ratio of the Bayes expected utility achieved by our proposed maximum likelihood rule to the Bayes expected utility achieved by each of the two alternative rules for even sample sizes between 2 and 200. The gains from using the optimal rule over either of the alternatives increase with the sample size. In a sample size of 200, the Bayes expected utility of the proposed maximum likelihood rule is 1.79 times higher than that of the Fréchet rule, and 1.56 times higher than that of the monotonicity rule.
Figure B.2: Under Optimality Conditions, For Increasing Sample Sizes, Performance of Proposed Maximum Likelihood Rule Increases Relative to Fréchet and Monotonicity Rules
Conditions for Bayes Optimality Enable the Construction of Credible Sets:
Under the conditions in which our proposed maximum likelihood decision rule is Bayes optimal, we can quantify uncertainty in estimation through Bayesian credible sets. We construct credible sets on the joint distribution of potential outcomes in the sample using the posterior belief implied by a uniform prior and the design-based likelihood. A credible set of level $\alpha$ is a set of joint distributions of potential outcomes covering a posterior probability mass of at least $\alpha$. We report the smallest credible set, which is the credible set containing the fewest points. We construct these sets with the following algorithm:
@@
We also report the range of values for each type—always takers, compliers, defiers, and never takers—present in the credible set constructed on the joint distribution. Note that the ranges reported on each type are not independent. For example, the maximal values of the ranges for always takers, compliers, defiers, and never takers may not constitute an element of the credible set on the joint distribution of potential outcomes.
Appendix B.4  Our Maximum Likelihood Decision Rule Maximizes Entropy
We can also motivate our maximum likelihood rule through the principle of maximum entropy (Jaynes, 1957b, a), which unites the theory of information with statistical physics. In Section 1.1, we discuss how differences in entropy drive variation in the likelihood such that the maximum likelihood distribution is also the maximum entropy distribution. Formally, the entropy is proportional to the logarithm of the number of assignments. Because the logarithm is monotonic, the distribution that maximizes the number of assignments also maximizes entropy. Jaynes (1968) advocates for the application of the principle of maximum entropy to circumvent the subjective specification of a prior. To make Bayesian decision-making more objective, one option is to choose the least informative prior. Jaynes’ alternative is to choose the least informative updated distribution. The distribution that can generate the data in the greatest number of ways—the distribution that maximizes entropy and our design-based likelihood—is the least informative updated distribution. In the Monty Hall game, a celebrated application of Bayesian decision making, it is possible to make the same decision without specifying a Bayesian prior using a maximum likelihood decision rule and the principle of maximum entropy (Wang, Tran, and Abebe, 2016).
Appendix C  Applications: Auxiliary Statistics
@@
Figure C.3: Standard Statistics, Proposed Design-Based Maximum Likelihood Estimates, and Auxiliary Statistics
Appendix D  Characterizing Rare Priors that Do Not Update About Monotonicity
Appendix D.1  Priors That Do Not Update About Monotonicity Can Have Undesirable Features
Since the release of an earlier version of this paper, Chen, Roth, and Spiess (2026) have demonstrated cases in which evidence revealed by the design-based likelihood has limited usefulness. These authors are primarily interested in evidence about a monotonicity hypothesis that there are no defiers, rather than evidence about the full distribution of potential outcomes. In particular, through a constructive proof, they show the existence of priors over the possible joint distributions of potential outcomes in the sample $\boldsymbol{\theta}$} under which the belief in a monotonicity hypothesis does not update for any realization of the data. Using our stylized example from Section 1.1, we demonstrate that priors constructed in this way can have undesirable features. First, priors $\pi$ constructed according to Chen, Roth, and Spiess (2026) for a sample of 6 people have limited support:
${\boldsymbol{\theta}}\in\text{supp }\pi\Rightarrow{\boldsymbol{\theta}}=\big(\theta_{11},\theta_{10},\theta_{01},\theta_{00}\big)\in\Big\{$ $(6,0,0,0),(4,0,0,2),(2,0,0,4),(0,0,0,6),$
$(0,1,5,0),(0,3,3,0),(0,5,1,0)\Big\}.$
There are ${6+4-1\choose 4-1}=84$ ways to allocate 6 people into 4 types, but these priors place weight on only 7—a violation of Cromwell’s rule, which advises against prior beliefs of 0 or 1 on any event (Lindley, 1985). Furthermore, these 7 consist only of distributions with two types present, so these priors place zero weight on any distribution with three or four types. Moreover, under such priors, the possible data realizations are only those in which an even number of subjects take up, so data with odd total takeup—like the data we consider in our stylized example, where 2 subjects take up in intervention and 1 takes up in control—are impossible under these priors. A researcher who adopted such a prior would have ruled out the data before seeing it.
Appendix D.2  Priors that Do Not Update About Monotonicity Are Rare
The undesirable features of the priors offered in the constructive proof of Chen, Roth, and Spiess (2026) motivate a characterization of priors that do not update about monotonicity. We consider the space of all priors that do not update about monotonicity, not just those constructed by Chen, Roth, and Spiess (2026). We show that priors that do not update about monotonicity are exceedingly rare, in the sense that they are both measure zero and nowhere dense.
We first establish some additional notation. Fix a finite sample size $n\geq 3$. Let $\boldsymbol{\Theta}$} be the set of joint distributions of types for sample size $n$. Define ${\boldsymbol{\Theta}}_{0}\equiv\{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}:\theta_{01}=0\ \text{or}\ \theta_{10}=0\}$ as the set of monotonic distributions and define ${\boldsymbol{\Theta}}_{1}$ as the set of non-monotonic distributions—the complement of ${\boldsymbol{\Theta}}_{0}$. Let $\Delta$ be the set of priors on $\boldsymbol{\Theta}$}:
$\Delta\equiv\bigg\{\pi:{\boldsymbol{\Theta}}\to[0,1]:\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\pi({\boldsymbol{\theta}})=1\bigg\}.$
Let $\mathcal{X}_{\pi}$ be the support of data realizations given a prior $\pi$, and define the set of non-updating priors $\mathcal{N}$:
$\mathcal{N}\equiv\{\pi\in\Delta:\pi({\boldsymbol{\Theta}}_{0}\mid{\boldsymbol{x}})=\pi({\boldsymbol{\Theta}}_{0})\,\forall{\boldsymbol{x}}\in\mathcal{X}_{\pi}\}.$
Proposition 2 presents our claim. We thank our excellent research assistant Shuheng Zhang for the formulation and proof of this proposition.
Proposition 2.
The set of non-updating priors $\mathcal{N}\subseteq\Delta$ is nowhere dense in $\Delta$ and has Lebesgue measure zero.
Proof.
For data realization $\boldsymbol{x}$}, define functions $A_{\boldsymbol{x}},B_{\boldsymbol{x}}:\Delta\to\mathbb{R}$ as:
$A_{\boldsymbol{x}}(\pi)\equiv\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}_{0}}\pi({\boldsymbol{\theta}})\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}}),\qquad B_{\boldsymbol{x}}(\pi)\equiv\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}_{1}}\pi({\boldsymbol{\theta}})\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}}).$
Define the function $F_{\boldsymbol{x}}:\Delta\to\mathbb{R}$ as:
$F_{\boldsymbol{x}}(\pi)=A_{\boldsymbol{x}}(\pi)\pi({\boldsymbol{\Theta}}_{1})-B_{\boldsymbol{x}}(\pi)\pi({\boldsymbol{\Theta}}_{0}).$
The posterior belief in monotonicity when ${\boldsymbol{x}}\in\mathcal{X}_{\pi}$ is given by $\pi({\boldsymbol{\Theta}}_{0}\mid{\boldsymbol{x}})=A_{\boldsymbol{x}}(\pi)(A_{\boldsymbol{x}}(\pi)+B_{\boldsymbol{x}}(\pi))^{-1}.$ A prior does not update its belief on monotonicity given a data realization $\boldsymbol{x}$} if and only if $A_{\boldsymbol{x}}(\pi)=\pi({\boldsymbol{\Theta}}_{0})(A_{\boldsymbol{x}}(\pi)+B_{\boldsymbol{x}}(\pi))$ by Bayes’ rule, or equivalently, if $F_{\boldsymbol{x}}(\pi)=0$. Note that $A_{\boldsymbol{x}}(\pi)$, $B_{\boldsymbol{x}}(\pi)$, $\pi({\boldsymbol{\Theta}}_{0})$, and $\pi({\boldsymbol{\Theta}}_{1})$ are all linear functions of the coordinates $\{\pi({\boldsymbol{\theta}})\}_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}$. Thus, $F_{\boldsymbol{x}}(\pi)$ is a quadratic polynomial in these coordinates. Now, for any given $\boldsymbol{x}$}, define $Z_{\boldsymbol{x}}$ as the set of all priors such that do not update their belief of monotonicity given $\boldsymbol{x}$}:
$Z_{\boldsymbol{x}}\equiv\big\{\pi\in\Delta:\pi({\boldsymbol{\Theta}}_{0}\mid{\boldsymbol{x}})=\pi({\boldsymbol{\Theta}}_{0})\big\}=\big\{\pi\in\Delta:F_{\boldsymbol{x}}(\pi)=0\big\}.$
The set $Z_{\boldsymbol{x}}$ is the intersection of $\Delta$ with the roots of the degree 2 polynomial $F_{\boldsymbol{x}}(\cdot)$.
Now, we will show there exists some ${\boldsymbol{x}}^{*}\in\mathcal{X}$ such that $F_{{\boldsymbol{x}}^{*}}(\cdot)$ is not the zero polynomial. Suppose for the sake of contradiction that $F_{\boldsymbol{x}}(\cdot)=0$ for all ${\boldsymbol{x}}\in\mathcal{X}$. Choose an $\boldsymbol{x}$}, ${\boldsymbol{\theta}}_{0}\in{\boldsymbol{\Theta}}_{0}$, and ${\boldsymbol{\theta}}_{1}\in{\boldsymbol{\Theta}}_{1}$. Consider the family of priors $\pi_{t}$ of the form $\pi_{t}({\boldsymbol{\theta}}_{0})=t$ and $\pi_{t}({\boldsymbol{\theta}}_{1})=1-t$ for $t\in(0,1)$, so $\pi_{t}({\boldsymbol{\theta}})=0$ for all other $\boldsymbol{\theta}$}. For such a $\pi_{t}$, we have that $A_{\boldsymbol{x}}(\pi_{t})=t\mathcal{L}({\boldsymbol{\theta}}_{0}\mid{\boldsymbol{x}})$, $B_{\boldsymbol{x}}(\pi_{t})=(1-t)\mathcal{L}({\boldsymbol{\theta}}_{1}\mid{\boldsymbol{x}})$, $\pi_{t}({\boldsymbol{\Theta}}_{0})=t$, and $\pi_{t}({\boldsymbol{\Theta}}_{1})=1-t$. Thus,
$F_{\boldsymbol{x}}(\pi_{t})$ $=t\mathcal{L}({\boldsymbol{\theta}}_{0}\mid{\boldsymbol{x}})(1-t)-(1-t)\mathcal{L}({\boldsymbol{\theta}}_{1}\mid{\boldsymbol{x}})t$
$=t(1-t)\big(\mathcal{L}({\boldsymbol{\theta}}_{0}\mid{\boldsymbol{x}})-\mathcal{L}({\boldsymbol{\theta}}_{1}\mid{\boldsymbol{x}})\big).$
Since $F_{\boldsymbol{x}}(\cdot)=0$, we have that $F_{\boldsymbol{x}}(\pi_{t})=0$ for all $t\in(0,1)$, which requires $\mathcal{L}({\boldsymbol{\theta}}_{0}\mid{\boldsymbol{x}})=\mathcal{L}({\boldsymbol{\theta}}_{1}\mid{\boldsymbol{x}})$. But $\boldsymbol{x}$}, ${\boldsymbol{\theta}}_{0}$, and ${\boldsymbol{\theta}}_{1}$ were arbitrarily chosen, which would imply that $\mathcal{L}({\boldsymbol{\theta}}_{0}\mid{\boldsymbol{x}})=\mathcal{L}({\boldsymbol{\theta}}_{1}\mid{\boldsymbol{x}})$ for all ${\boldsymbol{x}}\in\mathcal{X}$, ${\boldsymbol{\theta}}_{0}\in{\boldsymbol{\Theta}}_{0}$, and ${\boldsymbol{\theta}}_{1}\in{\boldsymbol{\Theta}}_{1}$, which contradicts our main likelihood curvature result. Thus, there exists at least one ${\boldsymbol{x}}^{*}\in\mathcal{X}$ such that $F_{{\boldsymbol{x}}^{*}}(\cdot)$ is not the zero function.
Fixing such a $\boldsymbol{x}$$^{*}$, we have that $F_{{\boldsymbol{x}}^{*}}(\pi)$ is polynomial and not the zero polynomial, so its roots form a set of measure zero. Thus, its intersection with $\Delta$, that is, the set $Z_{{\boldsymbol{x}}^{*}}$, is measure zero. Since
$\mathcal{N}=\bigcap_{{\boldsymbol{x}}\in\mathcal{X}}Z_{\boldsymbol{x}}\subset Z_{{\boldsymbol{x}}^{*}},$
it also must be measure zero. Then the set of priors that do not update on monotonicity is a set of measure zero. Furthermore, the interior of the closure of $Z_{{\boldsymbol{x}}^{*}}$ is empty, so $\mathcal{N}$ is nowhere dense. ∎
Proposition 2 implies that most Bayesians will learn about monotonicity, since the set of priors that do not update is measure zero. And, since this class of priors is nowhere dense in the space of priors, there are priors that cannot be closely approximated by non-updating priors. While Chen, Roth, and Spiess (2026) demonstrate the theoretical existence of priors that do not update about monotonicity, we show that these priors are rare.